The objective is to find the eigenvalues and corresponding eigenvectors. (2 0-1 1 0 Consider the matrix, A= 0 2 10 4

HOW TO FIND THE SECOND EIGEN VECTOR FOR A MULTIPLICITY 2 ?
Answer #1

а О Given Az 0 022 4 1 (A-2)=0 ão 0 72 0 0 07 1 0 Û O 4 0 ) -1 0 12-2 o zo. 1/2-2 o o 4-7 o. 2-9 [(4-) (4-1)] -1 (0-1 (1-21)-O 2-7 0 Ya-2 0 Y 4-2 2-12 13/2 d 9 1:03 o Z 7/2 Ri Rz - R3 3/2 0 -) 00 D o 2 Z 2574 25/4 Zz 20 Zzo 3% 2-Z=0 3/2 = 0 2 20 YzO0 -) 7-3 0 -5/2 0 7 0 R₂ + R₂ + R, To -1 O -) -5/2 0 cs 1-18) No 0 0 --720 -5), y = 0, Yz0 Zzt 2-1 -dat -a- zzo shoot d. o +t